The payroll tax for unemployment insurance in Patagonia taxes all wages up to a maximum per worker of 40,000 TL at a 10 percent flat rate. What are the marginal and average tax rates on the wages for each of the following three workers? a. A carpenter with annual wages of 38,000 TL b. An assistant manager with wages of 45,000 TL per year c. A corporate CEO with an annual salary of 600,000 TL
Answer #1

Marginal tax rate is 0% because tax is fixed at 10% for any income below 4000TL. Hence it is 0% for all a), b) and c)

Average tax rate for a) = 10% x 38000 / 38000 = 10%

Average tax rate for b) = 10% x 40000 / 45000 = 8.88%

Average tax rate of c) = 10% x 40000 / 600000 = 0.67%